# Automatically Generating UI Code from Screenshot: A Divide-and-Conquer-Based Approach

### v2

@@ 7.3. RQ3: How generalizable is DCGen for other SOTA MLLMs?
In this RQ, we examine the generalizability of DCGen by employing the methodology on different MLLMs as backbones. Tables 5 and 5 display the overall performance of DCGen on Gemini-1.5 and Claude-3. Our results indicate that DCGen is highly adaptable to the Gemini model, achieving notable gains of 3.9% in visual level metrics compared to other methods since in general, the decomposition allows models to conduct more focus on smaller visual segments and more manageable sub-generation tasks. For the Claude-3, DCGen also significantly enhances the visual similarity between the original and the generated websites by 3.8%, outperforming competing methods. As a result, the experiment demonstrates the generalizability of DCGen on various MLLMs, showing the superiority of the divide-and-conquer strategy on UI Code generation.
@@
Answer to RQ3: DCGen is effective across different MLLMs in generating UI code from design, which demonstrates the generalizability of the proposed divide-and-conquer strategy.
Table 4. DCGen’s performance on Gemini-1.5.
Gemini CLIP BLEU
Direct 72.01 8.46
CoT 72.10 8.65
Self-Refine 73.74 8.44
DCGen 77.63 8.81
Table 5. DCGen’s performance on Claude-3.
Claude-3 CLIP BLEU
Direct 73.77 5.94
CoT 73.21 5.99
Self-Refine 77.85 8.17
DCGen 81.68 6.00
@@
7.4. RQ4: What is the human’s feedback on DCGen in real-world application development workflows?
@@ 7.4.1. Pairwise Method Comparison
7.4.2. Usefulness Evaluation
Results. Figure 12 highlights UI implementation results from the experiment group (w/ DCGen) and control groups (w/o DCGen). First, the box plot in Figure 12(a) shows that the experiment group completes the design-to-code task twice as fast as the group without DCGen. The control group also shows greater variability, with some participants taking up to 20 minutes. In fact, the average time of the control group is underestimated, because one participant fail to complete at least one UI image within 20 minutes, which means that the participant needs more time in the real development. In contrast, all participants in the experiment group finish all the tasks within 15 minutes. Second, as shown in figure 12(b), the experiment group consistently achieves higher similarity scores compared to the control group. The Pearson Correlation Coefficient coefficient between the similarity scores given by the two judges in 12(b) is 0.71, suggesting a strong agreement between the judges. Finally, figure 12(c) summarizes these results, confirming that the experiment group using DCGen has lower completion times and higher similarity scores on average, indicating our tool can significantly help the experimental group implement webpages faster and more similar to the UI design images.
@@ 7.4.2. Usefulness Evaluation
Figure 12. The comparison of the experiment and control group on time and similarity.
Answer to RQ4: Human evaluations demonstrate that DCGen can produce more visually similar target webpages than competing methods and the DCGen tool can significantly help developers implement webpages faster and more similar to the UI designs.
@@ 8. Discussion
Figure 13. Success cases. The first row shows an example from RQ1, and the rest shows examples from the usefulness study (RQ4).
@@
Success stories. The first row of Figure 13 illustrates a success case from Section 7 RQ1, which shows that compared to the self-refine method, DCGen restores the color and layout of the webpage, and removes nonexistent pictures from the webpage thus reducing distortion. The rest of Figure 13 provides examples from the usefulness study (RQ4). DCGen demonstrates its effectiveness by significantly reducing element omissions and distortions, leading to faster development and improved webpage quality.
@@
In the usefulness study (RQ4), we observe that both developers in the control group follow a similar iterative refinement pattern to generate UI code, even though they are unaware of the DCGen’s methodology. Specifically, they start by prompting the model to produce an initial webpage, then iteratively refine the unsatisfactory parts by pointing out current errors and asking MLLMs for improvements in follow-up conversations. This process repeats until they are satisfied with the results or the time runs out. Meanwhile, developers also fix missing elements issues by providing more detailed prompts with a screenshot of the specific areas. Such manual narrowing of task scope aligns with DCGen’s automated divide-and-conquer approach.
@@
Tool features. Figure 14 shows a screenshot of DCGen’s user interface as a development tool, with the original image on the left and the code and rendered webpage on the right. DCGen enables a lot of features, allowing developers to generate an entire webpage (Figure 14(a)), inspect the code for individual image segments (Figure 14(b)), and regenerate code for any specific segment (Figure 14(c)). By automating the divide-and-conquer process, DCGen eliminates the need for detailed, repetitive prompts, speeding up development and enabling the creation of higher-quality webpages in less time compared to the control group.
(a) Generate code for the entire webpage.
(b) View the code of an image segment.
(c) Generate code for an image segment.
Figure 14. DCGen user interface and functionalities.
@@
Failure analysis. While DCGen effectively reduces element omission and distortion, its improvement in addressing element misarrangement is limited. A failure case is illustrated in Figure 14(a), within the orange rectangular area. In the original image, the paragraphs are horizontally arranged, but the generated paragraphs are vertically aligned. Upon further investigation, we found that this error occurs during the assembly process, as the leaf solver LLM correctly generates horizontally arranged paragraphs for this area, as shown in Figure 14(c). We attribute the issue to the inherent visual limitations of MLLMs, as identified by Tong et al. (Tong et al., 2024), which suggest that MLLMs still face challenges in recognizing certain simple visual patterns, such as positional arrangements.
@@
However, since DCGen allows users to review and regenerate code for any specific image segment, the erroneous code can be easily replaced with the correct code for the corresponding image segment in the final webpage. Thus, this failure case has minimal impact on DCGen’s overall practical effectiveness.
@@ 9. Threat to Validity
Limited context length. Current MLLMs have constraints on the length of the context window (e.g., 128K for GPT-4o). Consequently, our approach may fail to handle websites that require a vast number of tokens, exceeding the capacities of MLLMs. Nevertheless, our experiment demonstrates that all of the prompts is below the maximum context length, showing our approach is generalizable for most scenarios.
The selection of backbone models. This study employs three popular instruction-following and multimodal LLMs to demonstrate the effectiveness of DCGen. While other LLMs might be adapted for our methodology, we found that smaller parameter models fail to understand complex prompts. Our future work involves extending our implementation to other emerging models, showcasing DCGen’s generalization abilities.
@@ 10.1. Image-to-code Generation
The most informative AI techniques used for generating code from images can be classified into three broad categories: (i) Convolutional Neural Networks (CNNs) (Asiroglu et al., 2019; Beltramelli, 2018; Xu et al., 2021; Moran et al., 2018a; Chen et al., 2022c; Cizotto et al., 2023; Chen et al., 2018), (ii) Computer Vision (CV) and Optical Character Recognition (OCR) (Nguyen and Csallner, 2015; Natarajan and Csallner, 2018), and (iii) other deep learning models (Wu et al., 2021; Abdelhamid et al., 2020). The work (Beltramelli, 2018) utilizes CNNs and LSTM to extract features from GUI images to generate a domain-specific language (DSL). However, DSLs are not widely used in practical development, which inconveniences the industry’s UI developers (Xu et al., 2021). An alternative approach (Chen et al., 2018) using CNNs to extract visual features from UI images is more widely applicable than generating fixed DSL code. Based on CNNs, Andre et al. (Cizotto et al., 2023) pioneered using Class Activation Mapping (CAM), further improving the performance and interpretability of the image-to-code generation task.
@@ 10.2. Multimodal Language Models
Multimodal Language Models (MLLM) refer to LLM-based models capable of receiving, reasoning, and outputting multimodal information (Yin et al., 2023). In recent years, the trend of using LLMs as decoders in vision-language tasks has gained significant traction. This approach leverages cross-modal transfer, enabling knowledge sharing between language and multimodal domains. Pioneering studies such as VisualGPT (Chen et al., 2022b) and Frozen (Tsimpoukelli et al., 2021) have demonstrated the advantages of employing a pre-trained language model as a decoder. Following this, Flamingo (Alayrac et al., 2022) was developed to align a pre-trained vision encoder and language model using gated cross-attention, and it was trained on billions of image-text pairs, showcasing impressive in-context few-shot learning capabilities. Furthermore, BLIP-2 (Li et al., 2023b) introduced the use of Flan-T5 (Chung et al., 2024) with a Q-Former to efficiently align visual features with the language model. PaLM-E (Driess et al., 2023), featuring 562 billion parameters, was developed to integrate real-world continuous sensor modalities into an LLM, establishing a connection between real-world perceptions and human languages. Additionally, GPT-4 (OpenAI, [n. d.]) demonstrates enhanced visual understanding and reasoning abilities following pre-training on a vast collection of aligned image-text data. More recently, GPT-4o was proposed (OpenAI, 2024a) to achieve more effective and efficient image understanding and reasoning.

### v3

@@ 7.4. RQ4: How generalizable is DCGen for other SOTA MLLMs?
In this RQ, we examine the generalizability of DCGen by employing the methodology on different MLLMs as backbones. Tables 9 and 9 display the overall performance of DCGen on Gemini-1.5-flash and Gemini-2.0-flash. Our results indicate that DCGen is highly adaptable to the Gemini-1.5-flash model, achieving notable gains of 3.9% in visual level metrics. For Gemini-2.0, DCGen enhances the visual similarity between the original and the generated websites by 0.9%, outperforming competing methods. For Claude-3-Sonnet, DCGen also enhances the visual similarity by 3.8%, outperforming competing methods. As for code similarity, we observe that self-refine prompting can help Claude-3 preserve text content in the screenshots (e.g., paragraphs and headings), leading to a performance jump for code similarity under the self-refine prompt. Overall, the experiment demonstrates the generalizability of DCGen on various MLLMs, showing the superiority of the divide-and-conquer strategy on UI Code generation.
@@
Answer to RQ4: DCGen is effective across different MLLMs in generating UI code from design, which demonstrates the generalizability of the proposed divide-and-conquer strategy.
Table 7. Gemini-1.5-flash.
Gemini CLIP Code
Direct 72.0 16.6
CoT 72.1 16.8
Self-Refine 73.7 16.1
DCGen 77.6 17.8
Table 8. Gemini-2.0-flash.
Gemini CLIP Code
Direct 86.7 17.2
CoT 86.4 17.5
Self-Refine 86.5 17.5
DCGen 87.4 18.4
Table 9. Claude-3-Sonnet.
Claude-3 CLIP Code
Direct 73.8 13.5
CoT 73.2 13.3
Self-Refine 77.9 16.5
DCGen 81.7 13.5
@@
7.5. RQ5: What is the human’s feedback on DCGen in real-world application development workflows?
@@ 7.5.1. Pairwise Method Comparison
Figure 10. The comparison of the experiment (Exp) and control group (Ctrl) on time and similarity.
7.5.2. Usefulness Evaluation
Results. Figure 10 highlights UI implementation results from the experiment group (w/ DCGen) and control groups (w/o DCGen). First, the box plot in Figure 10(a) shows that the experiment group (Exp) completes the design-to-code task twice as fast as the group without DCGen. The control (Ctrl) group also shows greater variability, with some participants taking up to 20 minutes. The average time of the control group is underestimated because one participant failed to complete a UI image within 20 minutes. In contrast, all participants in the experiment group finish all the tasks within 15 minutes. Second, as shown in Figure 10(b), the experiment group consistently achieves higher similarity scores compared to the control group. The Pearson Correlation Coefficient between the similarity scores given by the two judges in Figure10(b) is 0.71, suggesting a strong agreement between the judges. Finally, Figure 10(c) summarizes these results, confirming that the experiment group using DCGen has lower completion times and higher similarity scores on average, indicating our tool can help the experimental group implement UI images faster and more accurately.
Answer to RQ5: Human evaluations demonstrate that DCGen can produce more visually similar target webpages than competing methods, and the DCGen tool can significantly help developers implement webpages faster and more similar to the UI designs.
@@ 8. Discussion
Figure 11. Success cases. The first row shows an example from RQ2, and the rest shows examples from the usefulness study (RQ5).
@@
Success stories. The first row of Figure 11 illustrates a success case from Section 7 RQ2, which shows that compared to the self-refine method, DCGen restores the color and layout of the webpage, and removes nonexistent pictures from the webpage thus reducing distortion. The rest of Figure 11 provides examples from the usefulness study (RQ5). DCGen demonstrates its effectiveness by significantly reducing element omissions and distortions, leading to faster development and improved webpage quality.
@@
In the usefulness study (RQ5), we observe that both developers in the control group follow a similar iterative refinement pattern to generate UI code, even though they are unaware of DCGen’s methodology. Specifically, they start by prompting the model to produce an initial webpage, then iteratively refine the unsatisfactory parts by providing more detailed prompts with a screenshot of the specific areas. Such manual narrowing of task scope aligns with DCGen’s automated divide-and-conquer approach.
@@
Tool features. Figure 12 shows a screenshot of DCGen’s user interface as a development tool, with the design image on the left and the code and rendered webpage on the right. DCGen enables developers to generate an entire webpage, inspect the code for individual image segments, and regenerate code for any specific segment. By automating the divide-and-conquer process, DCGen eliminates the need for detailed, repetitive prompts, speeding up development and enabling the creation of higher-quality webpages in less time compared to the control group.
Figure 12. Generate code for the entire webpage.
@@
Failure analysis. While DCGen effectively reduces element omission and distortion, its improvement in addressing element misarrangement is limited. A failure case is illustrated in Figure 12, within the orange rectangular area. In the original image, the paragraphs are horizontally arranged, but the generated paragraphs are vertically aligned. Upon further investigation, we found that this error occurs during the assembly process, as the leaf solver LLM correctly generates horizontally arranged paragraphs for this area, while the assembly LLM mistakenly changes the paragraphs to a vertical layout. We attribute the issue to the inherent visual limitations of MLLMs, as identified by Tong et al. (Tong et al., 2024), which suggest that MLLMs still face challenges in recognizing certain simple visual patterns, such as positional arrangements.
@@
However, since DCGen allows users to review and regenerate code for any specific image segment, the erroneous code segment can be easily replaced with the correct version in the final webpage. Thus, this failure case has minimal impact on DCGen’s practical effectiveness.
@@ 9. Threat to Validity
The selection of backbone models. This study employs four popular instruction-following and multimodal LLMs to demonstrate the effectiveness of DCGen. While other LLMs might be adapted for our methodology, we found that smaller parameter models fail to perform complex tasks such as the assembly process. Our future work involves extending our implementation to other emerging models, showcasing DCGen’s generalization abilities.
@@ 10.1. Image-to-code Generation
The most informative AI techniques used for generating code from images can be classified into three broad categories: (i) Convolutional Neural Networks (CNNs) (Asiroglu et al., 2019; Beltramelli, 2018; Xu et al., 2021; Moran et al., 2018a; Chen et al., 2022c; Cizotto et al., 2023; Chen et al., 2018), (ii) Computer Vision (CV) and Optical Character Recognition (OCR) (Nguyen and Csallner, 2015; Natarajan and Csallner, 2018), and (iii) other deep learning models (Wu et al., 2021; Abdelhamid et al., 2020). The work (Beltramelli, 2018) utilizes CNNs and LSTM to extract features from GUI images to generate a domain-specific language (DSL). However, DSLs are not widely used in practical development, which inconveniences the industry’s UI developers (Xu et al., 2021). An alternative approach (Chen et al., 2018) using CNNs to extract visual features from UI images is more widely applicable than generating fixed DSL code. Based on CNNs, Andre et al. (Cizotto et al., 2023) pioneered using Class Activation Mapping (CAM), further improving the performance and interpretability of the image-to-code generation task. More recently, Interaction2Code (Xiao et al., 2024) explored interactive web code generation from design images, and MRWeb (Wan et al., 2024) investigated the generation of multi-page resource-aware web code from UI designs, which extends page-level image-to-code generation to project-level.
@@ 10.2. Multimodal Language Models
Multimodal Language Models (MLLM) refer to LLM-based models capable of receiving, reasoning, and outputting multimodal information (Yin et al., 2023). In recent years, the trend of using LLMs as decoders in vision-language tasks has gained significant traction. Pioneering studies such as VisualGPT (Chen et al., 2022b) and Frozen (Tsimpoukelli et al., 2021) have demonstrated the advantages of employing a pre-trained language model as a decoder. Following this, Flamingo (Alayrac et al., 2022) was developed to align a pre-trained vision encoder and language model using gated cross-attention. Furthermore, BLIP-2 (Li et al., 2023b) introduced the use of Flan-T5 (Chung et al., 2024) with a Q-Former to efficiently align visual features with the language model. PaLM-E (Driess et al., 2023), featuring 562 billion parameters, was developed to integrate real-world continuous sensor modalities into an LLM. More recently, GPT-4o (OpenAI, 2024a) and Gemini-2.5 (Google, 2024) demonstrate enhanced visual understanding and reasoning abilities following pre-training on a vast collection of aligned image-text data.
@@ 11. Conclusion
Data Availability
All the code, data, and user-friendly tool for DCGen are available at https://github.com/WebPAI/DCGen for replication and future research.
Acknowledgment
This research was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant and Research Grants Council of the Hong Kong Special Administrative Region, China (No. SRFS2425-4S03 of the Senior Research Fellow Scheme).